OpenStudy (anonymous):

Given the graph of f'(x), the derivative of f(x), which of the following statements would be true about the graph of f(x)? I. The graph of f does not have a point of inflection. II. The graph of f has a point of inflection at x = 2. III. The graph of f has a critical point at x = 2.

OpenStudy (anonymous):

a. I only b. II only c. III only d. I and III only

OpenStudy (anonymous):

Please help, I really do need it.

OpenStudy (michele_laino):

please can you show us the graph of f'(x)?

OpenStudy (anonymous):

@ganeshie8 @Hero @AravindG @nincompoop @iGreen @agent0smith @johnweldon1993 @sammixboo @Preetha @AnswerMyQuestions @YanaSidlinskiy @Conqueror @Michele_Laino @linn99123

OpenStudy (michele_laino):

I think that at x=2 there is an inflection point for f(x)

OpenStudy (michele_laino):

since for x<2 f'(x) is an increasing function, and for x> 2 f'(x) is a decreasing function. So at x=2 there is a descendant inflection for f(x)

OpenStudy (anonymous):

Do you think any of the other answers are remotely correct?

OpenStudy (michele_laino):

I think the option b.

OpenStudy (anonymous):

Okay...I am still a little sketchy because I'm on my second try and if I get it wrong than I can't make that problem up. Thanks. Also, can you help me with this problem?

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

please is the graph of f(x) a parabola?

OpenStudy (anonymous):

here's all the information i have regarding the problem

OpenStudy (michele_laino):

I understand that tha graph of f(x) is a parabola. Now from that graph I inferred the subsequent equation for f(x): \[f(x)=-x ^{2}+9\] so we have: \[\frac{ d }{ dx }(-x ^{2}+9)=-2x\] and at x=3, I get f'(3)=-6

OpenStudy (anonymous):

thats not an answer tho

OpenStudy (michele_laino):

I think that your graph is not a parabola, so I think that: since the slope of the tangent line has to be negative at x=3, my answer is -1. the answer -5 can not be since the tangent line to your graph at x=3 is parallel to the bisector line of the II and IV quadrants

OpenStudy (anonymous):

can i see if we can get other opinions to clarify?

OpenStudy (anonymous):

@ganeshie8 @Hero @AravindG @nincompoop @iGreen @agent0smith @johnweldon1993 @sammixboo @Preetha @AnswerMyQuestions @YanaSidlinskiy @Conqueror @Michele_Laino @linn99123

OpenStudy (michele_laino):

ok!

OpenStudy (anonymous):

@ganeshie8 @hartnn @Hero @cwrw238 @iGreen @uri @Preetha

OpenStudy (anonymous):

Hey guys can you help I would really like a second opinion. Please?

OpenStudy (anonymous):

@mathmate Can you give us your opinion, please?

OpenStudy (anonymous):

It's the second problem

OpenStudy (mathmate):

@familyguymath Hey, I did some more checking. I have erred in the definition of inflection point... Sorry. An inflection point does not have to have f'(x)=0, even though it could be. As long as f"(x)=0, it is an inflection point, which means that as long as concavity switched, it is an inflection point. So you're right, (b) is the answer. I apologize for having spent much of your efforts, but we both benefit from learning in detail.

OpenStudy (anonymous):

Okay, thanks. But will you help us now on this new problem?

OpenStudy (anonymous):

@mathmate

OpenStudy (michele_laino):

I'm sorry but I think that the slope of the tangent line at x=3, can be -5 @familyguymath

OpenStudy (mathmate):

I agree with @Michele_Laino 's conclusions. We are looking at an estimate, which could be \(f'(x)=f(x)-f(x-1)\) which equals (9-4)=5. Michele has given a more exact answer by fitting the curve as a parabole over three points, (0,9),(1,8),(2,5),(3,0) which gives f(x)=9-x^2, and f'(x)=-2x, => f'(3)=-6 (as before).

OpenStudy (mathmate):

*parabola

OpenStudy (anonymous):

So you both think its the -5 and the -1

OpenStudy (anonymous):

because @Michele_Laino you put that a little while ago

OpenStudy (anonymous):

I calculated the slope as well and got and -6 but that # wasnt there so the closest # was the -5

OpenStudy (mathmate):

Estimates of slopes are better if it is symmetrical about the point inquestion, for example, f(3+0.5)-f(3-0.5)/(+0.5- -0.5). In this case, the graph ends at x=3, so we can more or less approximate.

OpenStudy (anonymous):

Hey guys but the regular slope is -5

OpenStudy (mathmate):

I don't know where -1 comes from, but -5 \(could\) be an answer obtained by approximating f'(3)=(f(3)-f(2))/1, graphically.

OpenStudy (anonymous):

ohh wait yah thats right

OpenStudy (mathmate):

@Michele_Laino :) I think you just confirmed graphically that the slope is -6. :)

OpenStudy (michele_laino):

wait, I'm sorry I have made an error in my drawing.

OpenStudy (anonymous):

I think that -5 is it? You guys agree?

OpenStudy (mathmate):

The question calls for an estimate, so anything from -5 to-7 \(should\) be acceptable.

OpenStudy (anonymous):

Will you guys check other problem?

OpenStudy (mathmate):

I agree that a good estimate is -6, but I can live with -5 because f(3)-f(2)=-5

OpenStudy (michele_laino):

here in my picture below, I have drawn the tangent line at graph at x=3, the bisector of the II and IV quadrants, and the line whose slope is -5

OpenStudy (mathmate):

Can you remind me of the "other problem"?

OpenStudy (anonymous):

For what values of x is the graph of y= (2)/5-x concave upward?

OpenStudy (anonymous):

i just want to know if x<5 was correct or not

OpenStudy (mathmate):

For all values of x where f"(x)>0. Is it \(\dfrac{2}{5-x}\) ?

OpenStudy (anonymous):

yes

OpenStudy (mathmate):

Well, f(x)=\(\dfrac{2}{5-x}\), f"(x)=\(\dfrac{4}{(5-x)^3}\) so x<5 will be all values x s.t. f"(x)>0. Yes, x<5 is correct.